OpenStudy (anonymous):

How do i find the missing coordinate when im given one coordinate and the midpoint. M=(-1,1) and P=(2,4)

OpenStudy (debbieg):

Use the midpoint formulas. The x-coordinate of the midpoint is given by:\[\Large M_x=\dfrac {x_{1}+x{2}}{2}\]So plug in \(\Large M_x\text{ and }{x_{1}}\), which you were given, and solve for \(\Large {x_{2}}\) Then do similarly for the y coordinate.

OpenStudy (anonymous):

im sorry that doesn't make sense to me :(

OpenStudy (debbieg):

Do you know the midpoint formula? If you have 2 points, do you know how to find their midpoint?

OpenStudy (anonymous):

yes

OpenStudy (debbieg):

\[\Large \left( \dfrac {x_{1}+x{_2}}{2}, \dfrac {y_{1}+y_{2}}{2} \right)\]

OpenStudy (anonymous):

yeah i know that formula but i dont understand how to use it to find the other point

OpenStudy (debbieg):

OK then, take just the x coordinate of the midpoint first. It is found by: \[\Large xmid=\dfrac {x_{1}+x{2}}{2}\] where i'm just using "xmid" to mean the x coordinate of the midpoint.

OpenStudy (debbieg):

Now, you're GIVEN the midpoint, so you know xmid, right? And you're GIVEN one point, so you know \(x_{1}\) because it's the x-coordinate of the known point. So plug those KNOWN values into that expression for xmid. then solve for \(x_{2}\)

OpenStudy (anonymous):

yeah i got -1= 2+x/2

OpenStudy (debbieg):

ok, good.... \[\Large -1=\dfrac {2+x}{2}\] Now solve that puppy for x!

OpenStudy (anonymous):

am i supposed to multiply -1 and 2

OpenStudy (debbieg):

Well, what should you do FIRST. take it one step at a time. How can you eliminate that fraction on the right side?

OpenStudy (anonymous):

Divide 2 and 2 so im left with -1=x?

OpenStudy (debbieg):

noooo.....

OpenStudy (debbieg):

if a=b/c Then ca=b

OpenStudy (anonymous):

IM SO LOST! :(

OpenStudy (debbieg):

So you clear a denominator of a fraction by MULTIPLYING by that denominator. Not dividing!

OpenStudy (anonymous):

so i should get -1=4+x

OpenStudy (debbieg):

Hang in there. I know you're frustrated but this is basic math and it's important that you work to wrap your head around it.

OpenStudy (anonymous):

does x=4?

OpenStudy (debbieg):

No. Two problems. 1. when you do ANYTHING on one side, you have to do it on the OTHER side too. You changed something on the right but didn't do anything on the left. 2. The numerator of the fraction is a whole term, that WHOLE THING is divided by 2. So when you multiply by 2, it cancels the den'r, but there's no "multiplying" in the numerator - it just cancels out \(\Large 2\cdot(-1)=\dfrac {2+x}{2}\cdot2\)

OpenStudy (debbieg):

And no, not x=4

OpenStudy (anonymous):

so im using the denominator when i multiply both sides?

OpenStudy (anonymous):

when i multiply the fraction do i multiply both twos by two?

OpenStudy (debbieg):

Here's an example with different expressions: \(\Large 3=\dfrac {1+x}{5}\) I want to clear that den'r of 5, so I multiply both sides by 5. \(\Large 5\cdot(3)=\dfrac {1+x}{5}\cdot5\) On right side, the five in den'r completely cancels with the 5 I'm multiplying by \(\Large 15={1+x}\)

OpenStudy (anonymous):

ok that makes more sense

OpenStudy (anonymous):

so right now i got -2=2+x

OpenStudy (debbieg):

no, here's why: \(\Large \dfrac {a}{b}\cdot b=a\) The WHOLE NUMERATOR is a.... whether it's a single number, a single variable, a sum or algebraic expression. The WHOLE NUMERATOR is one "unit" over the den'r, which cancels.

OpenStudy (debbieg):

GOOD!!

OpenStudy (debbieg):

Now you have -2=2+x So what's x? What do you do now?

OpenStudy (anonymous):

do i divide both sides by 2?

OpenStudy (debbieg):

No, no need to divide. You divide to undo a PRODUCT (multiplication). You don't have a product to deal with. You have a SUM (2+x). You need to get that 2 to the other side. What operation will "undo" a sum?

OpenStudy (anonymous):

subtraction

OpenStudy (debbieg):

Yes! Subtract 2 from both sides, to get the 2 to go "from" the right side "to" the left side. What do you get?

OpenStudy (anonymous):

x=0?

OpenStudy (debbieg):

Careful... you have a negative number on that left side. You are SUBTRACTING from a NEGATIVE. -2-2=?

OpenStudy (anonymous):

oh its -4 and for y i got -2.....is that correct? (-4,-2)

OpenStudy (debbieg):

YES!! Hooray!! :)

OpenStudy (anonymous):

alright thanks so much :)

OpenStudy (debbieg):

You're welcome. :) Good job sticking with it. Doesn't it feel good to have figured it out? :)

OpenStudy (anonymous):

yup :D